# LVLM-Count: Enhancing the Counting Ability of Large Vision-Language Models

### v3

@@ Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
LVLM-Count is a pipeline consisting of multiple stages. However, in the following section, we will demonstrate that the dominant portion of time during inference with a GPT-4o model is spent querying the LVLM to count objects, while other stages are negligible by comparison. Consequently, LVLM-Count can achieve inference times reasonably close to that of the base LVLM used.
In the four stages of the pipeline, the following main steps are taken: making a call to an LLM/LVLM to extract the object of interest from $Q$, calling GroundingDINO for area detection, calling GroundingDINO and SAM for target segmentation, performing some post-processing on the masks, running the mean-shift clustering algorithm, running the $A^{*}$ search, and querying the LVLM for the sub-images. From the listed steps, the running time of the post-processing on masks is negligible. The time required for the rest of the steps in an optimized pipeline is demonstrated in Table 12. To obtain the inference time for each step, we evaluated it on 1,000 images from the FSC-147 dataset and calculated the average. footnotetext: This is for a case where sub-image queries to the LVLM are made in parallel
Step Extract $E$ from $Q$ Area Detection Target Segmentation Clustering $A^{*}$ Sub-image Analysis Total Model Called GPT-4o GroundingDINO GroundingDINO SAM - - GPT-4o Time 0.56 0.10 0.10 0.33 0.24 0.03 1.65J 3.03
Table 12: Inference time breakdown of the LVLM-Count pipeline (Time in seconds)
As shown in Table 12, the dominant term in the inference time comes from calling GPT-4o to count objects of interest in the sub-images, which occupies more than half of the total inference time. However, a similar step occurs when querying the base LVLM to count objects of interest in the main image. The additional steps present in LVLM-Count but absent in the base LVLM constitute a smaller portion of the time. Table 13 presents a comparison between the total inference time of LVLM-Count and the base GPT-4o. This increase in inference time may be acceptable for most applications, given the substantial improvements in counting accuracy when using LVLM-Count.
@@
LVLM-Count (Using GPT-4o) 3.03
Table 13: Comparison of the inference time between the base GPT-4o and LVLM-Count
@@
Table 14 presents the results of dedicated counting models alongside LVLMs and their corresponding LVLM-Count variants on the FSC-147 test set. Note that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are the top three highest-performing text-based trained models on FSC-147 in the literature. Moreover, TFOC is the best-performing training-free model on FSC-147. We observe that, in general, the dedicated counting models (henceforth referred to simply as counting models) outperform the base LVLMs. However, the use of LVLM-Count significantly narrows this performance gap.
Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 24.79 - 137.15 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 14.90 - 103.42 - CountGD [Amini-Naieni et al., 2024] 14.76 - 120.42 - GroundingREC [Dai et al., 2024] 10.12 - 107.19 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
@@
Results on the PASCAL VOC dataset are reported in Table 15. For training-based counting models, this is considered a cross-dataset evaluation, meaning that they are trained on the training set of FSC-147 and tested on the PASCAL VOC benchmark. We observe that some trained counting models still outperform LVLMs; however, this is likely due to the significant overlap between the categories in PASCAL VOC and FSC-147. Nevertheless, LVLM-Count is very effective in narrowing the performance gap. Additionally, the results on the Emoji-Count benchmark are reported in Table 16. This is also a cross-dataset evaluation. In contrast to PASCAL VOC, Emoji-Count consists of complex concepts and categories that do not overlap with FSC-147. As a result, LVLMs, which have far stronger generalization power for unseen data and complex concepts, outperform counting models by a large margin. Employing LVLM-Count widens this gap even further.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 12.03 - 18.18 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 2.81 - 7.01 - GroundingRec [Dai et al., 2024] 4.05 - 7.80 - GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 198.99 - 208.08 - CountGD [Amini-Naieni et al., 2024] 137.93 - 156.80 - GroundingREC [Dai et al., 2024] 143.22 - 158.74 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 146.28$
@@
Finally, Table 17 presents the results of prior counting models on the Penguin benchmark alongside the counting performance of LVLMs. We observe that prior training-based counting models significantly outperform LVLMs on this benchmark. However, similar to other datasets, LVLM-Count narrows this performance gap.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 59.34 - 74.40 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 22.29 - 31.38 - CountGD [Amini-Naieni et al., 2024] 17.07 - 25.08 - GroundingREC [Dai et al., 2024] 22.04 - 26.75 - GPT4o 36.01 - 46.64 - LVLM-Count (GPT-4o as LVLM) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 Gemma 3 27B 49.87 - 61.06 - LVLM-Count (Gemma 3 27B as LVLM) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07
@@
Our focus in this work, similar to many recent works Liu et al. [2023], Amini-Naieni et al. [2023], is on text-based counting due to the above-mentioned reasons. Nevertheless, LVLM-Count is capable of working with visual exemplars alone as well as combinations of text prompts and visual exemplars, in addition to its text-only capabilities. The adaptation is straightforward, as most LVLMs also accept visual prompts. In Table 18, we evaluate the base GPT-4o and LVLM-Count (which uses GPT-4o) on Emoji-Count with text-only prompts, visual exemplars, and combinations of text and visual exemplars. For the visual-exemplar version, a single image of the target emoji icon is provided, and the LVLM is required to count similar instances. For the combined text-and-visual version, an image of the target emoji icon alongside its name is provided to the model. The results show higher accuracy when combining text and visual exemplars, while the text-only and exemplar-only versions achieve approximately the same performance.
@@
In this work, our primary focus is on developing a pipeline to enhance the counting ability of LVLMs for counting prompts in general, rather than optimizing for a specific prompting method. Nevertheless, we investigated the impact of Chain-of-Thought (CoT) prompting Wei et al. [2022] on visual counting task. For our experiments, we appended the phrase "think step-by-step" to the counting prompt. As shown in Table 19, the results on the PASCAL VOC benchmark reveal that CoT negatively affects counting performance—not only for LVLM-Count but also for the base LVLM.
@@
The zebra image is passed to the pipeline along with the question $Q=$“how many zebras are in the image?”. First, $E=$“zebra” is extracted using the LLM. Then the zebra image is passed to the area detection stage, where the prompt given to GroundingDINO is “zebras”. The output bounding boxes are merged, and the resulting area is cropped, as illustrated in Figure 21. The cropped area is then passed to the target segmentation stage. At this stage, GroundingDINO detects the objects of interest defined by $E$ as the input prompt. SAM then uses the output bounding boxes to produce segmentation masks for the zebras, as shown in Figure 22.
@@
Figure 22: Illustration of the target segmentation step of LVLM-Count for the zebra image. The goal is to produce all the instance masks for $E$ set to “zebra”. The cropped image from Figure 21, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 22(a). Figure 22(a) is then given as input to SAM, which produces the output shown in Figure 22(b).
@@
After the target segmentation stage, the masks are passed to the object-aware division stage. First, the masks are used in the cluster-based approach to find the location of the start and end points of the division paths, i.e., ($P^{1}_{s},P^{1}_{e}$) and ($P^{2}_{s},P^{2}_{e}$). Then these masks are turned into a black-and-white image, which, in turn, is mapped to a graph. The division paths are then found by connecting each start point to its corresponding end point by running the $A^{*}$ search algorithm on the graph. The found paths are mapped back into the image domain and drawn in red, as depicted in Figure 23. The image contours are determined based on the drawn red paths, and each contour’s interior is masked out independently to obtain the subimages. Finally, the subimages are given to the LVLM to count the number of zebras in each.

### v4

@@ Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
LVLM-Count is a pipeline consisting of multiple stages. However, in this section, we demonstrate that the dominant portion of the time during inference with the LVLMs tested in this paper, i.e. GPT4o, Gemma 3, and Qwen2, is spent querying the LVLM to count objects, and the additional steps in the pipeline comprise a small portion in comparison.
In the four stages of the pipeline, the following main steps are taken: making a call to an LLM/LVLM to extract the object of interest from $Q$, calling GroundingDINO for area detection, calling GroundingDINO and SAM for target segmentation, performing some post-processing on the masks, running the mean-shift clustering algorithm, running the $A^{*}$ search, and querying the LVLM for the sub-images. From the listed steps, the running time of the post-processing on masks is negligible. The time required for the rest of the steps in an optimized pipeline is demonstrated in Table˜12. To obtain the inference time for each step, we evaluated it on 1,000 images from the FSC-147 dataset and calculated the average. footnotetext: This is for a case where sub-image queries to the LVLM are made in parallel
Table 12: Inference time breakdown of the LVLM-Count pipeline for different LVLMs (Time in seconds)
Method Extract $E$ from $Q$ Area Detection Target Segmentation Clustering $A^{*}$ Subimage AnalysisJ Total LVLM-Count (GPT4o) 0.11 0.10 0.43 0.24 0.03 1.65 2.56 LVLM-Count (Gemma 3 27B) 0.11 1.11 2.02 LVLM-Count (Qwen2 VL 72B AWQ) 0.10 1.17 2.07
As shown in Table˜12, the dominant term in the inference time comes from calling the LVLM to count objects of interest in the sub-images. This occupies more than half of the total inference time. However, a similar step occurs when querying the base LVLM to count objects of interest in the main image. The additional steps present in LVLM-Count but absent in the base LVLM constitute a smaller portion of the time. Table˜13 presents a comparison between the total inference time of LVLM-Count and the base LVLMs. This increase in inference time may be acceptable for most applications, given the substantial improvements in counting accuracy when using LVLM-Count.
Table 13: Comparison of the inference time between the base LVLM and LVLM-Count
@@
LVLM-Count (Using GPT-4o) 2.56
Base Gemma 3 27B 1.52
LVLM-Count (Using Gemma 3 27B) 2.01
Base Qwen2 VL 72B AWQ 1.46
LVLM-Count (Using Qwen2 VL 72B AWQ) 2.07
@@
Table˜14 presents the results of dedicated counting models alongside LVLMs and their corresponding LVLM-Count variants on the FSC-147 test set. Note that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are the top three highest-performing text-based trained models on FSC-147 in the literature. Moreover, TFOC is the best-performing training-free model on FSC-147. We observe that, in general, the dedicated counting models (henceforth referred to simply as counting models) outperform the base LVLMs. However, the use of LVLM-Count significantly narrows this performance gap.
@@
Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 24.79 - 137.15 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 14.90 - 103.42 - CountGD [Amini-Naieni et al., 2024] 14.76 - 120.42 - GroundingREC [Dai et al., 2024] 10.12 - 107.19 - GPT-4o 25.57 - 137.26 - LVLM-Count (GPT-4o as LVLM) 17.86 $\downarrow 7.71$ 91.71 $\downarrow 45.55$ Gemma 3 27B 30.59 - 132.61 - LVLM-Count (Gemma 3 27B as LVLM) 20.25 $\downarrow 10.34$ 110.45 $\downarrow 22.16$ Qwen2 VL 72B AWQ 34.48 - 149.49 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.29 $\downarrow 11.89$ 119.46 $\downarrow 30.44$
Results on the PASCAL VOC dataset are reported in Table˜15. For training-based counting models, this is considered a cross-dataset evaluation, meaning that they are trained on the training set of FSC-147 and tested on the PASCAL VOC benchmark. We observe that some trained counting models still outperform LVLMs; however, this is likely due to the significant overlap between the categories in PASCAL VOC and FSC-147. Nevertheless, LVLM-Count is very effective in narrowing the performance gap. Additionally, the results on the Emoji-Count benchmark are reported in Table˜16. This is also a cross-dataset evaluation. In contrast to PASCAL VOC, Emoji-Count consists of complex concepts and categories that do not overlap with FSC-147. As a result, LVLMs, which have far stronger generalization power for unseen data and complex concepts, outperform counting models by a large margin. Employing LVLM-Count widens this gap even further.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 12.03 - 18.18 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 2.81 - 7.01 - GroundingRec [Dai et al., 2024] 4.05 - 7.80 - GPT4o 4.64 - 8.56 - LVLM-Count (GPT4o as LVLM) 3.42 $\downarrow$ 1.22 7.17 $\downarrow$ 1.39 Gemma 3 27B 3.40 - 7.54 - LVLM-Count (Gemma 3 27B as LVLM) 2.97 $\downarrow$ 0.0.43 6.16 $\downarrow$ 1.38 Qwen2 VL 72B AWQ 4.80 - 8.71 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.12 $\downarrow$ 0.68 7.76 $\downarrow$ 0.95
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 198.99 - 208.08 - CountGD [Amini-Naieni et al., 2024] 137.93 - 156.80 - GroundingREC [Dai et al., 2024] 143.22 - 158.74 - GPT-4o 23.57 - 36.97 - LVLM-Count (GPT-4o as LVLM) 16.57 $\downarrow 7$ 33.11 $\downarrow 3.86$ Gemma 3 27B 21.39 - 24.04 - LVLM-Count (Gemma3 27B as LVLM) 16.16 $\downarrow 5.23$ 21.27 $\downarrow 2.77$ Qwen2 VL 72B AWQ 78.05 - 159.22 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 24.43 $\downarrow 53.62$ 43.38 $\downarrow 115.84$
Finally, Table˜17 presents the results of prior counting models on the Penguin benchmark alongside the counting performance of LVLMs. We observe that prior training-based counting models significantly outperform LVLMs on this benchmark. However, similar to other datasets, LVLM-Count narrows this performance gap.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 59.34 - 74.40 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 22.29 - 31.38 - CountGD [Amini-Naieni et al., 2024] 17.07 - 25.08 - GroundingREC [Dai et al., 2024] 22.04 - 26.75 - GPT4o 35.18 - 45.76 - LVLM-Count (GPT-4o as LVLM) 26.76 $\downarrow$ 8.42 38.60 $\downarrow$ 7.16 Gemma 3 27B 49.44 - 60.20 - LVLM-Count (Gemma 3 27B as LVLM) 34.13 $\downarrow$ 15.31 44.11 $\downarrow$ 16.09 Qwen2 VL 72B AWQ 44.02 - 65.59 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.21 $\downarrow$ 15.81 40.40 $\downarrow$ 25.19
@@
Our focus in this work, similar to many recent works Liu et al. [2023], Amini-Naieni et al. [2023], is on text-based counting due to the above-mentioned reasons. Nevertheless, LVLM-Count is capable of working with visual exemplars alone as well as combinations of text prompts and visual exemplars, in addition to its text-only capabilities. The adaptation is straightforward, as most LVLMs also accept visual prompts. In Table˜18, we evaluate the base GPT-4o and LVLM-Count (which uses GPT-4o) on Emoji-Count with text-only prompts, visual exemplars, and combinations of text and visual exemplars. For the visual-exemplar version, a single image of the target emoji icon is provided, and the LVLM is required to count similar instances. For the combined text-and-visual version, an image of the target emoji icon alongside its name is provided to the model. The results show higher accuracy when combining text and visual exemplars, while the text-only and exemplar-only versions achieve approximately the same performance.
@@
In this work, our primary focus is on developing a pipeline to enhance the counting ability of LVLMs for counting prompts in general, rather than optimizing for a specific prompting method. Nevertheless, we investigated the impact of Chain-of-Thought (CoT) prompting Wei et al. [2022] on visual counting task. For our experiments, we appended the phrase "think step-by-step" to the counting prompt. As shown in Table˜19, the results on the PASCAL VOC benchmark reveal that CoT negatively affects counting performance—not only for LVLM-Count but also for the base LVLM.
@@
The zebra image is passed to the pipeline along with the question Q=Q=“how many zebras are in the image?”. First, E=E=“zebra” is extracted using the LLM. Then the zebra image is passed to the area detection stage, where the prompt given to GroundingDINO is “zebras”. The output bounding boxes are merged, and the resulting area is cropped, as illustrated in Figure˜21. The cropped area is then passed to the target segmentation stage. At this stage, GroundingDINO detects the objects of interest defined by $E$ as the input prompt. SAM then uses the output bounding boxes to produce segmentation masks for the zebras, as shown in Figure˜22.
@@
Figure 22: Illustration of the target segmentation step of LVLM-Count for the zebra image. The goal is to produce all the instance masks for $E$ set to “zebra”. The cropped image from Figure˜21, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure˜22(a). Figure˜22(a) is then given as input to SAM, which produces the output shown in Figure˜22(b).
@@
After the target segmentation stage, the masks are passed to the object-aware division stage. First, the masks are used in the cluster-based approach to find the location of the start and end points of the division paths, i.e., ($P^{1}_{s},P^{1}_{e}$) and ($P^{2}_{s},P^{2}_{e}$). Then these masks are turned into a black-and-white image, which, in turn, is mapped to a graph. The division paths are then found by connecting each start point to its corresponding end point by running the $A^{*}$ search algorithm on the graph. The found paths are mapped back into the image domain and drawn in red, as depicted in Figure˜23. The image contours are determined based on the drawn red paths, and each contour’s interior is masked out independently to obtain the subimages. Finally, the subimages are given to the LVLM to count the number of zebras in each.